# Finitely many worlds with finite information flow and Bell theorem

### v1

@@ Abstract
The Copenhagen interpretation of quantum theory does not provide a measurement-independent description of physical processes. Two competing approaches to this issue are single-world ontological theories and Everett many-worlds theory. The former addresses the measurement problem by introducing auxiliary random variables that specify, in particular, the actual values of macroscopic observables. Such models may be $\psi$-epistemic, meaning the quantum state is not part of the ontology. However, all the ontological models break the principle of locality, as demonstrated by Bell’s theorem. The latter approach, which is deterministic and $\psi$-ontic, offers the advantage of evading Bell’s theorem by delaying the buildup of the correlations until the parties compare their outcomes at a meeting point. Because of determinism and the ontic nature of the quantum state, the number of parallel worlds and the information flow towards the meeting point turn out to be infinite. By integrating the strengths of the two approaches, we introduce a simple $\psi$-epistemic, many-worlds, local model of projective measurements on two spatially separate maximally entangled qubits. Because of its randomness, the model requires only two branches and a finite information flow – just one bit per measurement is communicated to the meeting point. We explore how this hybrid approach, employing both randomness and branching, addresses key challenges of each individual framework.
.0.1 Introduction
@@
Since the inception of quantum theory, interpretative challenges have raised questions about its completeness epr . To reconcile state superpositions with the definiteness of the our macroscopic experience, a minimal requirement is the existence of auxiliary information that determines the actual macroscopic state that is experienced. This idea is pursued in ‘hidden-variable’ theories and is consistently realized in the de Broglie-Bohm (dBB) theory bohm . In the dBB theory, both the wave-function and the position of the particles specify the actual state of a system. Employing a term coined in the context of quantum foundations, the state represented by the wave-function and the auxiliary variables is referred as an ontic state. More broadly, instead of variables, the precise ontic state itself could be hidden. Therefore, more generally, we refer to these hypothetical completions as ontological theories.
It is important to emphasize that ontological theories are inherently statistical, reflecting the randomness of quantum phenomena. In Bohmian theory, for instance, particle positions are randomly distributed according to Born’s rule. As we will see, however, the randomness of a theory of quantum phenomena is not self-evident.
@@
A possible resolution of this issue is to assume that there is a multitude of branches such that the ratio between the branches in |+ket|$\rangle$| + ⟩ and those in |ket|$\rangle$| - ⟩ is about $|\psi_{+}/\psi_{-}|^{2}$ saunders . Since this ratio is a real number, the Born rule is recovered exactly in the limit of infinite branches. Thus, if a system splits into a finite number of macroscopically distinct states, each one actually corresponds to an infinite multitude of parallel worlds. Hereafter, in the context of many-worlds theory, we employ this derivation of Born rule.
While (single-world) ontological theories are probabilistic, the many-worlds theory is $\psi$-ontic and deterministic. Randomness arises from the subjective perspective of an observer following a single branch of the tree. In this paper, we explore how integrating the randomness of ontological theories with many-worlds branching can yield a more economical framework while preserving an advantage of many-worlds theory – its claimed ability to provide a local, realistic picture of the world(s) deutsch . The argument supporting this ability is clearly illustrated by a simple local realistic model of a ‘nonlocal’ Popescu-Rohrlich (PR) box brassard . The core idea of the argument is to delay the buildup of correlations until the parties compare their outcomes at a meeting point. However, many-worlds theory requires an infinite number of branching and an infinite information flow to the meeting point deutsch .
In this paper, we show that just one classical bit of information about the chosen measurement and two branches are sufficient to simulate local projective measurements on maximally entangled qubits. This drastic reduction of information is a consequence of the probabilistic and $\psi$-epistemic nature of the proposed model. While in the many-worlds theory quantum states are part of the ontology, which must therefore contain infinite classical information, in our model the quantum state is rather encoded in the statistics of the ontic states.
@@
.0.2 Nonlocal correlations in many-worlds theory
@@
$|\Psi\rangle=\frac{1}{\sqrt{2}}\left(|-1,-1\rangle-|1,1\rangle\right).$
The qubits are sent to two spatially separated parties, Alice and Bob. Each party independently chooses to perform a projective measurement, represented by the Bloch vectors $\vec{a}$ and $\vec{b}$, respectively. The outcomes, ±$\pm 1$, of each measurement occur with equal probability. In the MW framework, this means that qubits locally branch into a multitude of realities in half of which a party observes one of the two outcomes. On each side, this process occurs without the need to know what is happening on the other side. After the measurements, the parties decide to compare their results. To do that, they have to meet at the same place or send some communication through a classical channel. After multiple running of the experiment over different entangled states, they find out that the outcomes, $s_{A}\in\{-1,1\}$ and $a_{B}\in\{-1,1\}$, are distributed according to the probability distribution
@@
which violate the Bell inequalities. While this violation leads to a break of local realism in a single-world scenario, the MW framework evades this conclusion as follows. When Alice meets with Bob, the multitudes of realities of each party are paired in such a way that the fraction of realities experiencing the outcomes $s_{A}$ and $s_{B}$ is exactly equal to $P(s_{A},s_{B}|\vec{a},\vec{b})$. This argument is illustrated in Ref. brassard for nonlocal PR boxes and further discussed in the general quantum case in Ref. charles . The existence of many local realities allows us to delay the buildup of the correlations until the parties meet or communicate each other.
It is fundamental to remark that, in the MW theory, each party has to carry the information about the performed measurement at the meeting point in order to get the right pairing. This information is clearly infinite, which is a consequence of the $\psi$-ontic nature of the theory.
.0.3 Finitely many worlds with finite information flow
Now, let us show that two branches and a finite information flow are enough for simulating the outcomes of all the projective measurements $\vec{a}$ and $\vec{b}$ on a maximally entangled state. The proof is a straightful application of MW argument to Toner-Bacon model toner . The model provides a classical simulation of the outcomes by using one bit of communication between the parties and some shared randomness.
@@
Now, let us reinterpret this model within a many-worlds framework, where each party branches into two distinct alternatives, labeled $A_{\pm 1}$ and $B_{\pm 1}$ for Alice and Bob, respectively. First, we note that the Toner-Bacon model generates the same correlation if $s_{A}$ and $s_{B}$ are both chosen with opposite sign. Let us assume that alternatives $A_{1}$ and $A_{-1}$ get the outcome $s_{A}$ in Eq. (3) and its opposite, respectively. They both generate $n_{A}$ according to Eq. (4). All these processes involve only local information that is available to Alice. A complication arises in deciding which outcomes are associated with Bob’s alternatives. Associating $s_{B}$ in Eq. (5) to alternative $B_{1}$ would require to know the value $n_{A}$, an information that is not available to Bob. If we set $n_{A}$ equal to $1$ and, finally, we pair alternative $A_{w}$ with $B_{w}$ at the meeting point, we end up making a mistake in the case $n_{A}$ were $-1$ and $\text{sign}\left[\vec{b}\cdot\vec{x}_{+}\right]\neq\text{sign}\left[\vec{b}\cdot\vec{x}_{-}\right]$, where $\vec{x}_{\pm}\equiv\vec{x}_{0}\pm\vec{x}_{1}$. Crucially, the mistake can be corrected by knowing just one classical bit of information per measurement, namely $n_{A}$ for Alice’s measurement and
@@
This model is mathematically equivalent to the model in Ref. gisin , but they differ in the interpretation. The latter is a single-world model in which the parties share a nonlocal PR box as a resource for generating the outcome. In our model, we have a ”local” PR box which is used for pairing locally the alternatives at the meeting point. The statistics of the single-world model of Ref. gisin is recovered by randomly taking one of the two paired alternatives in the ”many-worlds” model.
@@
.0.4 Discussion
The nonlocality arising in the framework of ontological theories can lead the proponents of many-worlds theory to claim the failure of a probabilistic framework. The main purpose of this paper is to show with a proof of principle that only the branching of the worlds is required for a local account of the correlations of two maximally entangled qubits. There is no known proof that both determinism and the ontology of the quantum states are necessary conditions for a local theory. Furthermore, the drop of these two assumptions can lead to a more ‘economical’ theory, in which the information flow and the number of branches is made as small as possible.
@@
The second difference is more subtle. In a $\psi$-epistemic theory, the ontic state does not contain the full information about the quantum state. Rather, this information is encoded in the statistics of the ontic state. Let us consider the quantum state emerging from Alice’s measurement and how its information is encoded in many-worlds theory and our model. In the former, the ongoing state, which has to be carried to the meeting point, is the vector $\vec{a}$ and the outcome associated with each branch. Their knowledge allows us to infer the outgoing quantum state after the measurement in each branch, meaning that the quantum state is part of the ontology. In the latter, this inference is not possible in a single run of the experiment. After the measurement, Alice generates a number $n_{A}$ and associates an outcome to each branch. This information and the knowledge of the shared random variables $\vec{x}_{0}$ and $\vec{x}_{1}$ enable us only to infer that the quantum state $\vec{a}$ in branch $A_{1}$ after the measurement is any unit vector such that Eqs. (3,4) are fulfilled. Rather, the full information about quantum state is statistically encoded in the ontic variables. This is necessary for making the information flow from the parties to the meeting point finite.
.0.5 Conclusions
@@
In conclusion, a hybrid approach that combines the randomness of ontological theories with the branching structure of many-worlds theory can address the challenges of each individual framework. On one hand, branching allows us to evade Bell’s theorem, which applies to single-world ontological theories. More broadly, it may also circumvent the contextuality requirement. For instance, a manifestation of preparation contextuality is the proof that ontological theories inherently break time symmetry leifer1 . However, with appropriate adaptations, our model provides a fully time-symmetric description of a scenario in which a qubit, in a maximally mixed state, is measured at two different times. The time asymmetry proved in Ref. leifer1 is transferred to the measurement devices and the subsequent comparison of results, which are inherently time-asymmetric processes. Similarly, the branching circumvents the recently proved theorem that measurements erase information monti_erase , which is another manifestation of contextuality. On the other hand, the introduction of randomness in the many-worlds framework can make it more natural and economical to account for the observed randomness by allowing us to reduce the number of branches and the information flow as much as possible – it may be possible that there is no infinite branching and not all the alternatives in a superposition are necessarily realized. If there were some branch with a very faint amplitude in which Scipio was defeated in the Third Punic War, this would not necessarily imply that there is a world in which Carthage rules over Rome!
@@
Here, we have presented a proof of principle of this hybridation by showing that local measurements on two maximally entangled qubits can be locally simulated with only two branches, finite communication and shared randomness. It is plausible that a generalization to many qubits and any quantum state will require a blow-up of the shared variables. This increase of local resources occurs also in the local version of the many-worlds theory deutsch , in which the local observables are represented by matrices acting on the whole Hilbert space of all involved particles. The main difference, however, is that the shared random variables in our approach would not depend on the state of the system – they would be generated before the ”game” starts and would provide a stochastic background over which the system evolves. A similar idea has been suggested in Ref. montina_non_markov within a single-world scenario. However, whether this hybrid approach can be extended to a general theory remains an open and, in our opinion, fascinating question.
.0.6 Acknowledgments

### v2

# A local, many-worlds, model of quantum correlations with finite information flow

@@ Abstract
Ontological theories, such as the de Broglie-Bohm theory, address the measurement problem by introducing auxiliary random variables that specify, in particular, the actual values of macroscopic observables. Such models may be $\psi$-epistemic, meaning the quantum state is not part of the ontology. A serious issue of this route toward a realistic completion of quantum theory is raised by Bell’s proof that ontological theories are nonlocal. A possible resolution is to reject the assumption that measurements have single actual outcomes. Indeed, relaxing this premise, Deutsch and Hayden showed that Bell’s theorem can be evaded by delaying the buildup of the correlations until the parties compare their outcomes at a meeting point. However, the Deutsch-Hayden theory, which is determinist and $\psi$-ontic, leads to an infinite information flow towards the meeting point. Furthermore, alternative branches are weighted by amplitudes, leading to interpretative issues. By integrating the randomness of single-world theories and the branching of the Deutsch-Hayden theory, we introduce a simple $\psi$-epistemic local model of projective measurements on two spatially separate maximally entangled qubits. Because of its randomness, the model requires two ”equally weighted” branches and a finite information flow – just one bit per measurement is communicated to the meeting point. We explore how this hybrid approach, employing both randomness and branching, addresses key challenges of single-world and Deutsch-Hayden theories. On one hand, the branching allows us to circumvent nonlocality and, possibly, contextuality. On the other hand, randomness makes it more natural and economical to derive quantum probabilities from unweighted counts of branches and ensemble averages. Furthermore, it allows for a reduction of the information flow by stripping the quantum state of its ‘ontic’ rank.
I Introduction
@@
Since the inception of quantum theory, interpretative challenges have raised questions about its completeness epr . To reconcile state superpositions with the definiteness of the our macroscopic experience, a minimal requirement is the existence of auxiliary information that determines the actual macroscopic state that is experienced. This idea is pursued in ‘hidden-variable’ theories and is consistently realized in the de Broglie-Bohm (dBB) theory bohm . In the dBB theory, both the wave-function and the position of the particles specify the actual state of a system. Employing a term coined in the context of quantum foundations, the state represented by the wave-function and the auxiliary variables is referred as an ontic state. More broadly, instead of some variables, the precise ontic state itself could be hidden. Therefore, more generally, we refer to these hypothetical completions as ontological theories.
It is important to emphasize that ontological theories are inherently probabilistic, reflecting the randomness of quantum phenomena. In Bohmian theory, for instance, particle positions are randomly distributed according to Born’s rule. As we will see, however, the randomness of a theory of quantum phenomena is not self-evident.
@@
A pragmatic approach is to assume that our expectation of being in a particular branch is given Born’s rule, which is adopted as a postulate. A concrete derivation of the rule from a branch counting is obtained by assuming that there is a multitude of branches such that the ratio between the branches in |+ket|$\rangle$| + ⟩ and those in |ket|$\rangle$| - ⟩ is about $|\psi_{+}/\psi_{-}|^{2}$, as proposed by Saunders saunders . Since this ratio is a real number, the Born rule is recovered exactly in the limit of infinite branches. Thus, if a system splits into a finite number of macroscopically distinct states, each one actually corresponds to an infinite multitude of parallel worlds. Hereafter, in the context of many-worlds theory, we just assume that branches are weighted by an amplitude and, for the sake of concreteness, we also refer to the branch counting of Ref. saunders .
While (single-world) ontological theories are probabilistic, the many-worlds theory is $\psi$-ontic and deterministic. Randomness arises from the subjective perspective of an observer following a single branch of the tree. In this paper, we explore how integrating the randomness of ontological theories with many-worlds branching can yield a more economical framework while preserving an advantage of Deutsch-Hayden (DH) many-worlds theory deutsch – its claimed ability to provide a local, realistic picture of the world(s). The argument supporting this ability is clearly illustrated by a simple local realistic model of a ‘nonlocal’ Popescu-Rohrlich (PR) box brassard . The core idea of the argument is to delay the buildup of correlations until the parties compare their outcomes at a meeting point. However, DH theory requires branch weights and an infinite information flow to the meeting point deutsch . Employing the branch counting of Ref. saunders , also the number of branches turns out to be infinite.
In this paper, by adopting a hybrid framework that incorporates both randomness and branching, we show that just one classical bit of information about the chosen measurement and two “equally weighted” branches are sufficient to simulate local projective measurements on maximally entangled qubits. Randomness allows us to derive the quantum probabilities more naturally and economically from a simple unweighted count of branches and an ensemble average. Furthermore, the $\psi$-epistemic nature of the model leads to a drastic reduction of the information flow. While in the many-worlds theory quantum states are part of the ontology, which must therefore contain infinite classical information, in our model the quantum state is rather encoded in the statistics of the ontic states.
@@
II Nonlocal correlations in many-worlds theory
@@
$|\Psi\rangle=\frac{1}{\sqrt{2}}\left(|-1,1\rangle-|1,-1\rangle\right).$
The qubits are sent to two spatially separated parties, Alice and Bob. Each party independently chooses to perform a projective measurement, represented by the Bloch vectors $\vec{a}$ and $\vec{b}$, respectively. The outcomes, ±$\pm 1$, of each measurement occur with equal probability. In the DH framework, this means that qubits locally split into two branches with equal weights. Rephrasing in Saunders’ terms, there is a multitude of realities in half of which a party observes one of the two outcomes. On each side, this process occurs without the need to know what is happening on the other side. After the measurements, the parties decide to compare their results. To do that, they have to meet at the same place or send some communication through a classical channel. After multiple running of the experiment over different entangled states, they find out that the outcomes, $s_{A}\in\{-1,1\}$ and $a_{B}\in\{-1,1\}$, are distributed according to the probability distribution
@@
which violate the Bell inequalities. While this violation leads to a break of local realism in a single-world scenario, the DH framework evades this conclusion as follows. When Alice meets with Bob, four branches emerge such that each one is weighted with an amplitude whose modulus square is exactly $P(s_{A},s_{B}|\vec{a},\vec{b})$. Crucially, these weights are not in conflict with the previous weight assignments. Rephrasing in Saunders’ terms, the multitudes of realities of each party are paired in such a way that the fraction of realities experiencing the outcomes $s_{A}$ and $s_{B}$ is equal to $P(s_{A},s_{B}|\vec{a},\vec{b})$. This argument is illustrated in Ref. brassard for nonlocal PR boxes and further discussed in the general quantum case in Ref. charles . The existence of many local realities allows us to delay the buildup of the correlations until the parties meet or communicate each other.
It is fundamental to remark that, in the DH theory, each party has to carry the information about the performed measurement at the meeting point in order to get the right pairing. This information is clearly infinite, which is a consequence of the $\psi$-ontic nature of the theory.
III A model with finite information flow
Now, let us show that two “equally weighted” branches and a finite information flow are enough for simulating the outcomes of all the projective measurements $\vec{a}$ and $\vec{b}$ on a maximally entangled state. The proof is a straightful application of DH argument to Toner-Bacon model toner . The model provides a classical simulation of the outcomes by using one bit of communication between the parties and some shared randomness.
@@ III A model with finite information flow
Now, let us reinterpret this model within a framework à la many-worlds, where each party branches into two distinct alternatives, labeled $A_{\pm 1}$ and $B_{\pm 1}$ for Alice and Bob, respectively. The two parties are not allowed to communicate until they reach the meeting point. First, we note that the Toner-Bacon model generates the same correlation if $s_{A}$ and $s_{B}$ are both chosen with opposite sign. Let us assume that alternatives $A_{1}$ and $A_{-1}$ get the outcome $s_{A}$ in Eq. (3) and its opposite, respectively. They both generate $n_{A}$ according to Eq. (4). All these processes involve only local information that is available to Alice. A complication arises in deciding which outcomes are associated with Bob’s alternatives. Associating $s_{B}$ in Eq. (5) to alternative $B_{1}$ would require to know the value $n_{A}$, an information that is not available to Bob. If we set $n_{A}$ equal to $1$ and, finally, we pair alternative $A_{w}$ with $B_{w}$ at the meeting point, we end up making a mistake in the case $n_{A}$ were $-1$ and $\text{sign}\left[\vec{b}\cdot\vec{x}_{+}\right]\neq\text{sign}\left[\vec{b}\cdot\vec{x}_{-}\right]$, where $\vec{x}_{\pm}\equiv\vec{x}_{0}\pm\vec{x}_{1}$. Crucially, the mistake can be corrected by knowing just one classical bit of information per measurement, namely $n_{A}$ for Alice’s measurement and
@@
From the perspective of one realization of Alice and Bob, the probability of getting outcomes $s_{A}$ and $s_{B}$ is obtained by a simple “unweighted” branch counting. First, we compute the fraction of branches getting $s_{A}$ and $s_{B}$. Then, we average over $\vec{x}_{0}$ and $\vec{x}_{1}$. Note that there is no amplitude weight of the branches.
This model is mathematically equivalent to the model in Ref. gisin , but they differ in the interpretation. The latter is a single-world model in which the parties share a nonlocal PR box as a resource for generating the outcome. In our model, we have a ”local” PR box which is used for pairing locally the alternatives at the meeting point. The statistics of the single-world model of Ref. gisin is recovered by randomly taking one of the two paired alternatives in the ”many-worlds” model. It is surprising that a simple and seemingly trivial conceptual step separates our model from that in Ref. gisin , yet it has gone unnoticed until now.
@@
IV Discussion
The nonlocality arising in the framework of ontological theories can lead the proponents of Deutsch-Hayden many-worlds theory to claim the failure of a probabilistic framework. The main purpose of this paper is to show with a proof of principle that only the branching is required for a local account of the correlations of two maximally entangled qubits. There is no known proof that both determinism and the ontology of the quantum states are necessary conditions for a local theory. Furthermore, the drop of these two assumptions can lead to a more ‘economical’ theory, in which the information flow and the number of branches is made as small as possible.
Let us discuss in detail the two key differences between our model and the Deutsch-Hayden many-worlds theory – namely, the way the probabilities come out and the role played by the quantum state.
Probabilities – In the scenario simulated by the model, if Alice and Bob perform the measurements $\vec{a}$ and $\vec{b}$, then they observe outcome $s_{A}$ and $s_{B}$ with probability $(1-s_{A}s_{B}\vec{a}\cdot\vec{b})/2$. The many-worlds theory says that there are $4$ branches, one for each outcome, weighted by an amplitude whose modulus square is the observed probability. In Saunders’ terms, there is a multitude of branches such that the ratio between the number of branches with outcomes $s_{A}$ and $s_{B}$ and the number of overall branches is equal to the observed probability. Since, this probability is a real number, this multitude is actually infinite.
In our model, the story is quite different. There are only two “equally weighted” branches and, in a single run of the experiment, the number of branches in which an outcome is observed can be 00 or $1$. This implies that some outcomes are not realized in a single instance of the experiment (in MW theory, all the outcomes are realized). The observed probability of an outome is obtained by taking the fraction of branches with that outcome and averaging over different runs. This allows us to recover Born’s rule in a more natural and economical way.
@@ IV Discussion
Role of the quantum state – The second difference is more subtle. In a $\psi$-epistemic theory, the ontic state does not contain the full information about the quantum state. Rather, this information is encoded in the statistics of the ontic state. Let us consider the quantum state emerging from Alice’s measurement and how its information is encoded in Deutsch-Hayden theory and our model. In the former, the ongoing state, which has to be carried to the meeting point, is the vector $\vec{a}$ and the outcome associated with each branch. Their knowledge allows us to infer the outgoing quantum state after the measurement in each branch, meaning that the quantum state is part of the ontology. In the latter, this inference is not possible in a single run of the experiment. After the measurement, Alice generates a number $n_{A}$ and associates an outcome to each branch. This information and the knowledge of the shared random variables $\vec{x}_{0}$ and $\vec{x}_{1}$ enable us only to infer that the quantum state $\vec{a}$ in branch $A_{1}$ after the measurement is any unit vector such that Eqs. (3,4) are fulfilled. Rather, the full information about quantum state is statistically encoded in the ontic variables. This is necessary for making the information flow from the parties to the meeting point finite montina_comm_epist0 ; montina_comm_epist .
V Conclusions
@@
In conclusion, a hybrid approach that combines the randomness of ontological theories with the branching structure of many-worlds theory can address the challenges of each individual framework. On one hand, branching allows us to evade Bell’s theorem, which applies to single-world ontological theories. More broadly, it may also circumvent the contextuality requirement. For instance, a manifestation of preparation contextuality is the proof that ontological theories inherently break time symmetry leifer1 . However, with appropriate adaptations, our model provides a fully time-symmetric description of a scenario in which a qubit, in a maximally mixed state, is measured at two different times (this model will be presented in a future version of Ref. monti_erase ). The time asymmetry proved in Ref. leifer1 is transferred to the measurement devices and the subsequent comparison of results, which are inherently time-asymmetric processes. Similarly, the branching circumvents the recently proved theorem that measurements erase information monti_erase , which is another manifestation of contextuality. A many-worlds framework can also elude the Pusey-Barrett-Rudolph theorem pbr , which implies that locality is in conflict with $\psi$-epistemic theories under a single-world premise. Note that this bypass of the theorem has never been considered so far, since Deutsch-Hayden theory is local but also $\psi$-ontic. On the other hand, the introduction of randomness in the many-worlds framework can make it more natural and economical to account for the observed randomness through a simple count over a finite set of branches. Furthermore, it allows us to reduce the information flow as much as possible by stripping the quantum state of its ‘ontic’ rank. Finally, if the quantum state represents our knowledge about a system rather than an element of reality, then it is possible that not all alternatives in a superposition are necessarily realized (as occurs in our model). If there were some branch with a very faint amplitude in which Scipio was defeated in the Third Punic War, this would not necessarily imply that there is a world in which Carthage rules over Rome!
@@
Here, we have presented a proof of principle of this hybridation by showing that local measurements on two maximally entangled qubits can be locally simulated with only two ‘unweighted’ branches, finite communication and shared randomness. It is plausible that a generalization to many qubits and any quantum state will require a blow-up of the shared variables. This increase of local resources occurs also in the Deutsch-Hayden many-worlds theory deutsch , in which the local observables are represented by matrices acting on the whole Hilbert space of all involved particles. The main difference, however, is that the shared random variables in our approach would not depend on the state of the system – they would be generated before the ”game” starts and would provide a stochastic background over which the system evolves. A similar idea has been suggested in Ref. montina_non_markov within a single-world scenario. However, whether this hybrid approach can be extended to a general theory remains an open and, in our opinion, fascinating question.
Acknowledgments